Find sin(theta), sec(theta), and cot(theta) if tan(theta)=2/7
Answer by lwsshak3(11628)

Find sin(theta), sec(theta), and cot(theta) if tan(theta)=2/7
***
use x for theta
assume reference angle x is in quadrant I
tanx=2/7
hypotenuse of reference right triangle=√(2^2+7^2)=√4+49=√53
sinx=2/√53=2√53/53
cosx=7/√53=7√53/53
secx=√53/7
cotx=7/2